Consider a relation R=\ (1,1),(1, ), (0,2), (2,3) (3,1)) on the set A=\ 1,2,3\ Find transitive closure of the relation R using algorithm Warshall's

1
Expert's answer
2021-12-21T18:31:50-0500

Let us state the steps of the Warshall's algorithm:


Step 1. Let "W:=M_R,\ k:=0."

Step 2. Put "k:=k+1."

Step 3. For all "i\ne k" such that "w_{ik}=1" and for all "j" let "w_{ij}=w_{ij}\lor w_{kj}."

Step 4. If "k=n" then STOP: "W=M_{R^*}," else go to Step 2.


Consider a relation "R=\{ (1,1),(1, 0), (0,2), (2,3), (3,1)\}" on the set "A=\{0, 1,2,3\}".


Let us find transitive closure of the relation "R" using Warshall's algorithm:


"W^{(0)}=M_R
=\begin{pmatrix}
0 & 0 & 1 & 0\\
1 & 1 & 0 & 0\\
0 & 0 & 0 & 1\\
0 & 1 & 0 & 0
\end{pmatrix}"



"W^{(1)}
=\begin{pmatrix}
0 & 0 & 1 & 0\\
1 & 1 & 1 & 0\\
0 & 0 & 0 & 1\\
0 & 1 & 0 & 0
\end{pmatrix}"


"W^{(2)}
=\begin{pmatrix}
0 & 0 & 1 & 0\\
1 & 1 & 1 & 0\\
0 & 0 & 0 & 1\\
1 & 1 & 1 & 0
\end{pmatrix}"


"W^{(3)}
=\begin{pmatrix}
0 & 0 & 1 & 1\\
1 & 1 & 1 & 1\\
0 & 0 & 0 & 1\\
1 & 1 & 1 & 1
\end{pmatrix}"


"M_{R^*}=W^{(4)}
=\begin{pmatrix}
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1\\
1 & 1 & 1 & 1
\end{pmatrix}"


It follows that "R^*=A\times A" is a universal relation on the set "A."